OpenStudy (clara1223):

Determine if the Mean Value Theorem applies to the function f(x)=2sin(x)+sin(2x) on the interval [0,pi]. If so, find all numbers c on the interval that satisfy the theorem.

zepdrix (zepdrix):

Which one is the Mean Value Theorem again? Oh right right right.. the one with the uhhhh... You have some interval... With the `secant line` connecting the end points... Mean Value Theorem tells us that there is a `tangent line` somewhere inside of that interval. So ummm

OpenStudy (clara1223):

yeah so you have to find (f(pi)-f(0))/(pi-0)

OpenStudy (clara1223):

and then set that equal to the derivative of the function

OpenStudy (clara1223):

i just have trouble with trig on this kind of problem

ganeshie8 (ganeshie8):

how do you know MVT can be applied here? have you shown that the given function meets the requirements necessary for applying MVT ? 

OpenStudy (clara1223):

I graphed it and it is continuous and differentiable on the interval

ganeshie8 (ganeshie8):

Fair enough! now you're ready to apply MVT

ganeshie8 (ganeshie8):

As a start, find the slope of secant line of f(x)=2sin(x)+sin(2x) between [0,pi]

ganeshie8 (ganeshie8):

f(x)=2sin(x)+sin(2x) f(pi) = 0 f(0) = 0 so, clearly (f(pi)-f(0))/(pi-0) = 0

ganeshie8 (ganeshie8):

so, you want to solve \[2\cos(c)+\cos(2c)=0\] over the interval \((0,\pi)\)

ganeshie8 (ganeshie8):

recall the identity : \(\cos(2x)=2\cos^2x-1\)

OpenStudy (clara1223):

and how do i use that identity?

ganeshie8 (ganeshie8):

\(2\cos(c)+2\cos(2c)=0\) \(\cos(c)+\cos(2c)=0\) Applying double angle identity : \(\cos(c)+2\cos^2(c)-1=0\) \(2\cos^2(c)+\cos(c)-1=0\) Try factoring it